for each quadriatic function(a) find the vertex, the axis of symmetry, anf the maximum or minimum function value and (b) graph the function.
f(x) = 4Xsquare + 16x +13

Answer by Alan3354(69443)

a) find the vertex, the axis of symmetry, anf the maximum or minimum function value
The axis of symmetry is x = -b/2a
x = -16/8 = -2
x = -2 is the axis
----------
The minimum is at the vertex which is on the axis of symmetry
f(-2) = 16 -32 + 13 = -3
Minimum at (-2,-3)
Vertex at (-2,-3)
----------------
and (b) graph the function.
f(x) = 4x^2 + 16x +13
dl the FREE graph software at
http://www.padowan.dk.com